# Assessing the Adversarial Security of Practical Perceptual Hashing Algorithms

### v1

@@ III-B2 Robustness to Adversarial Blackbox Attacks
Proposition 1. Gradient estimation based on queried hash is not reliable.
@@
Proof. Without loss of generality, consider the optimization problem (4). A common approach involves utilizing $K$ randomly perturbed images to query the PHAs and then using the queried hashes to estimate the gradient as follows
@@
III-C Hash Inversion Attacks
GAN (generative adversarial network) was used by [12]. However, the service provider adversary may need more efficient hash inversion models because the training data may not be abundant in practice. For this, we train new and more efficient generative models to reconstruct the original input images from the hashes. We designed a decider style DNN, similar to the Fast Style Transfer [32], to convert the $N$-bit hashes to $(64,64)$ images. This DNN is mainly a combination of residual blocks, which are described in Fig. 1(b), with fractionally-strided convolutions/deconvolutions that are used for learned upsampling. The entire architecture is described in Fig. 1(a).
For a given training dataset, we first compute the hashes of each image offline and store the image-hash pairs. We then train the model outlined above to reconstruct the images as closely as possible by minimizing the mean-square-error (MSE) loss between the DNN output and the image from which the hash was generated.
IV Experiment
IV-A Performance Metrics
For assessing the robustness against blackbox attacks, we employ the attack success rate as the evaluation metric. Given a set of $M$ original images ${\bf I}_{0}^{m}$ with their corresponding hashes ${\bf h}_{0}^{m}$, the adversary generates adversarial images ${\bf I}_{\rm adv}^{m}$ along with their hashes ${\bf h}_{\rm adv}^{m}$. The attack success rate is defined as
${\rm ASR}(p)=$
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left(D_{h}({\bf h}_{0}^{m},{\bf h}_{\rm adv}^{m})>p)\right)\mathbb{I}\left(L_{2}({\bf I}_{0}^{m},{\bf I}_{\rm adv}^{m})<\theta\right)$ (7)
for untargeted attacks, and
${\rm ASR}(p)=$
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left(D_{h}({\bf h}_{\rm tar}^{m},{\bf h}_{\rm adv}^{m})<p)\right)\mathbb{I}\left(L_{2}({\bf I}_{0}^{m},{\bf I}_{\rm adv}^{m})<\theta\right)$ (8)
for targeted attacks (with the target hash ${\bf h}_{\rm tar}$). $p\in[0,1]$ is the pre-set hash distance threshold and $\theta$ is the pre-set image distortion threshold. The adversary can adjust the threshold $p$ to control the tradeoff between the true positive rate (TPR) and the false positive rate (FPR) of the hash or image similarity decisions.
Untargeted attacks use a large $p$ because ${\bf h}_{\rm adv}$ should be different from ${\bf h}_{0}$. As shown in Fig. 2, when we calculated $D_{h}({\bf h}_{0},{\bf h}_{1})$ of distinct images of the ImageNet validation dataset, we found that very few pairs of images had $D_{h}({\bf h}_{0},{\bf h}_{1})<0.3$, which was also observed in [11]. Therefore, we primarily use $p=0.3$ in (7) to determine the success of untargeted attacks. In contrast, targeted attacks use a small $p$ to make sure ${\bf h}_{\rm adv}$ is sufficiently close to ${\bf h}_{\rm tar}$. Prior studies [11] indicate that many normal image edits would change an image’s hashes to the level of $D_{h}({\bf h}_{0},{\bf h}_{1})<0.1$. Therefore, we mainly use $p=0.1$ in (8) for assessing the success of targeted attacks.
To comprehensively assess attack success, we collect ASR$(p)$ for all $p=0.1$, $0.2$, $0.3$, $0.4$ during experiments. We set the image distortion threshold $\theta=0.1$.
Figure 2: Cumulative distribution function (cdf) of normalized hamming distance $D_{h}({\bf h}_{0},{\bf h}_{1})$ between hashes of distinct images in the ImageNet validation dataset.
For hash inversion attacks, we use the $L_{2}$ distortion (2), the SSIM (Structural Similarity Index Measure) [33], and the LPIPS (Learned Perceptual Image Patch Similarity) [34] between the reconstructed image and the true image to evaluate the performance. No ASR or query budget is involved.
IV-B Experiment Results of Image Editing Attacks
@@
Experiment results are presented in Tables I. The data for ASR(.3) is highlighted in black because $p=0.3$ is the threshold we suggested to focus on. It is evident that all three PHAs were robust to compression and resizing, but not to filtering and rotation. In conjunction with the findings in [11][25], the PHAs are robust to many standard image editing operations except the following:
@@
NeuralHash was the most robust among the three. Surprisingly, the newest edition of NeuralHash that we experimented with was still not robust to filtering, rotating, and mirroring, similar to the first edition experimented in [11][25], even though DNNs can be trained to be invariant to such standard operations with data augmentation.
TABLE I: Adversary’s ASR($p$) with Image Editing Attacks
@@
IV-C Experiment Results of Untargeted Blackbox Attacks
In implementing our proposed JSHA algorithm, we modified three different soft-label attack algorithms: SimBA [21], NES [19], and ZOsignSGD [20], to use in Step One, and modified the hard-label attack algorithm HSJA [22] to use in Step Two. For instance, “SimBA+HSJA” means our JSHA with a modified SimBA as Step One and a modified HSJA as Step Two. Modifications included removing/changing constraints and adjusting hyper-parameters, etc. Specifically, in Step One, we skipped image distortion constraints and adopted extremely large perturbation step sizes, e.g, using a perturbation magnitude of $\epsilon=0.9$ for SimBA instead of the original $\epsilon=0.01$ as described in [21].
Experiment results are presented in Tables II. We compared the proposed JSHA with the straightforward application of standard SimBA [21], NES [19], and ZOsignSGG [20] algorithms. We also compared it with [26] and [28].
With a focus on ASR(0.3), the data indicated that NeuralHash was relatively robust because the maximum ASR$(0.3)$ was $28\%$ only. PDQ was robust against all adversarial attacks, with a maximum ASR$(0.3)$ of 1% only. It’s worth noting that although [26] achieved a much higher attack success rate against PDQ, it required significantly more queries. It could afford millions of queries due to its access to the PDQ source code, enabling faster computation.
Furthermore, the experiment data revealed that our proposed JSHA outperformed standard blackbox attack algorithms. SimBA+HSJA surpassed [28] in attacking PhotoDNA, as the latter produced more noisy adversarial images and utilized more queries.
Sample adversarial images generated with our proposed JSHA with various pre-set threshold $p$ are depicted in Fig. 3, with their corresponding distortion metrics listed in Table III. Notably, the three images with $p=0.4$ were deemed invalid adversarial images as their distortion was not under the threshold $\theta=0.1$, resulting in extremely noisy images. Similarly, for PDQ, the adversarial images with $p=0.2$ and $0.3$ were also invalid.
In summary, contrary to previous studies suggesting that PHAs were not robust against untargeted attacks, we found that some PHAs like PDQ and NeuralHash could be considered as robust if realistic constraints regarding image distortion and query budget are taken into account.
TABLE II: Adversary’s ASR($p$) with Untargeted Black-box Attacks.
(PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
@@ IV-C Experiment Results of Untargeted Blackbox Attacks
(a) Original image ${\bf I}_{0}$.
PhotoDNA                  PDQ                   NeuralHash
(b) Adversarial images ${\bf I}_{\rm adv}$.
TABLE III: Distortion of the Adversarial Image Samples in Fig. 3(b).
@@
IV-D Experiment Results of Targeted Blackbox Attacks
The experiment setting of targeted attacks was similar to that of untargeted attacks, except for utilizing ${\bf h}_{\rm tar}$ obtained from random images instead of ${\bf h}_{0}$. Results are presented in Tables IV. With a focus on ASR(0.1) in this case, we observed that all targeted attacks failed, yielding ASR$(0.1)=0$. To our knowledge, no literature has ever reported non-zero ASR$(0.1)$, suggesting that all the PHAs are robust against targeted attacks.
The robustness is extremely strong because of ASR$(p)\approx 0$ across all $p$. The only exception was NeuralHash, which exhibited ASR(0.4) of 43% with SimBA. However, this isn’t indicative of successful adversarial attacks; rather, it’s due to numerous hashes generated by NeuralHash having $D_{h}({\bf h}_{0},{\bf h}_{1})$ as low as 0.3, as illustrated in Fig. 2. Under NeuralHash, many images are inherently adversarial when evaluated at $p=0.4$.
In summary, contrary to the prior claims that PHAs were not robust against targeted attacks [28], our findings demonstrate the strong robustness of all PHAs against targeted blackbox attacks. This is significant to the practical application of PHAs because the client adversary needs targeted attacks to ensure that its adversarial hashes do not inadvertently match illicit hashes in the database. The strong robustness means it is hard for the adversary to undertake either untargeted or targeted attacks without being caught.
TABLE IV: Adversary’s ASR($p$) with Targeted Blackbox Attacks.
(PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
ZO- ZOsign
SimBA NES+ sign- -SGD
ASR SimBA +HSJA NES HSJA SGD +HSJA
PH ASR(.1) 0% 0% 0% 0% 0% 0%
PH ASR(.2) 0% 0% 0% 0% 0% 0%
PH ASR(.3) 0% 0% 0% 0% 0% 0%
PH ASR(.4) 1% 2% 0% 0% 0% 0%
PD ASR(.1) 0% 0% 0% 0% 0% 0%
PD ASR(.2) 0% 0% 0% 0% 0% 0%
PD ASR(.3) 0% 0% 0% 0% 0% 0%
PD ASR(.4) 0% 0% 0% 0% 0% 0%
NH ASR(.1) 0% 0% 0% 0% 0% 0%
NH ASR(.2) 0% 0% 0% 0% 0% 0%
NH ASR(.3) 0% 0% 0% 0% 0% 0%
NH ASR(.4) 43% 11% 0% 0% 0% 0%
[28] PH ASR(.3) ($L_{2}=0.19$) 83% ($2\times 10^{6}$ queries)
[28] PD ASR(.3) ($L_{2}=0.42$) 100% ($6\times 10^{5}$ queries)

### v2

@@ 3.3 A Defense Method based on Random Hash Perturbation
Proposition 1. Gradient estimation based on queried hash is not reliable due to the random variation of hash bits.
An outline of the proof is in the Appendix A.1. The random variation of hash bits makes the adversaries’ optimization struggle to converge, which will be demonstrated in the next section. Furthermore, considering this observation, we propose a defense method where the PHAs randomly invert $q$-portion of the hash bits to guarantee sufficient hash randomness, where $0\leq q\leq 1$.
4 Experiment
4.1 Performance Metrics
We use the attack success rate for assessing the robustness against blackbox adversarial hash-evasion attacks. Given a set of $M$ original images ${\bf I}_{0}^{m}$ with their corresponding hashes ${\bf h}_{0}^{m}$, the adversary generates adversarial images ${\bf I}_{\rm adv}^{m}$ with hashes ${\bf h}_{\rm adv}^{m}$. The attack success rate is defined as
${\rm ASR}(p)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left(D_{h}({\bf h}_{0}^{m},{\bf h}_{\rm adv}^{m})\gtrless p)\right)\mathbb{I}\left(L_{2}({\bf I}_{0}^{m},{\bf I}_{\rm adv}^{m})<\theta\right)$ (3)
[…]
For the thresholds, we set $\theta=0.1$. On the other hand, we should apply large enough $p$ to evaluate untargeted attacks because ${\bf h}_{\rm adv}$ should be different from ${\bf h}_{0}$. It was found in [11] that most pairs of distinct images had $D_{h}({\bf h}_{0},{\bf h}_{1})>0.3$. Therefore, we primarily use $p=0.3$ in (3) to determine the success of untargeted attacks. In contrast, we should use small enough $p$ for targeted attacks to make sure ${\bf h}_{\rm adv}$ is sufficiently close to ${\bf h}_{\rm tar}$. Prior studies [11] indicate that many normal image edits would change an image’s hashes to the level of $D_{h}({\bf h}_{0},{\bf h}_{1})<0.1$. Therefore, we mainly use $p=0.1$ in (3) for assessing the success of targeted attacks. This important parameter $p$ was unfortunately neglected by most existing adversarial robustness studies.
For hash-inversion attacks, we use the $L_{2}$ distortion, the SSIM (Structural Similarity Index Measure) [26], and the LPIPS (Learned Perceptual Image Patch Similarity) [27] between the reconstructed image and the true image to evaluate the performance. No ASR or query budget is involved.
4.2 Experiment on Hash-Evasion Attack and Defense
In implementing our proposed JSHA algorithm, we modified three different soft-label attack algorithms: SimBA [28], NES [29], and ZOsignSGD [30], to use in Step One, and modified the hard-label attack algorithm HSJA [31] to use in Step Two. For instance, “SimBA+HSJA” means our JSHA with a modified SimBA as Step One and a modified HSJA as Step Two. Modifications included removing/changing constraints and adjusting hyper-parameters, etc. We compared the proposed JSHA with the direct application of the original SimBA, NES, and ZOsignSGD algorithms. We also compared it with existing works [17] and [19].
For targeted hash-evasion attacks, experiment results are presented in Table 1 and Appendix A.3. With a focus on ASR(0.1), we observed that all targeted attacks failed, yielding ASR$(0.1)=0$. To our knowledge, no literature has ever reported non-zero ASR$(0.1)$, suggesting that all the PHAs are robust against blackbox targeted hash-evasion attacks.
Table 1: ASR($p$) of Targeted Hash-Evasion Attacks. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
ZO- ZOsign
SimBA NES+ sign- -SGD
ASR SimBA +HSJA NES HSJA SGD +HSJA
PH ASR(.1) 0% 0% 0% 0% 0% 0%
PD ASR(.1) 0% 0% 0% 0% 0% 0%
NH ASR(.1) 0% 0% 0% 0% 0% 0%
[19] PH ASR(.3) ($L_{2}=0.19$) 83% ($2\times 10^{6}$ queries)
[19] PD ASR(.3) ($L_{2}=0.42$) 100% ($6\times 10^{5}$ queries)
For untargeted attacks, experiment results are presented in Table 2 and Appendix A.4. With a focus on ASR(0.3), we see that PDQ was robust against all adversarial attacks, with a maximum ASR$(0.3)$ of 1% only. It’s worth noting that although [17] achieved a much higher attack success rate against PDQ, it required significantly more queries. The results also demonstrated that our defense method was effective because it drastically reduced ASR(0.3) and made the PHAs robust.
Table 2: Adversary’s ASR($p$) in Untargeted Blackbox Hash-Evasion Attacks, either without defense ($q=0$) or with defense $q=0.1$. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
ZO- ZOsign Defense Defense
SimBA NES+ sign- -SGD +SimBA +NES
ASR SimBA +HSJA NES HSJA SGD +HSJA +HSJA +HSJA
PH ASR(.3) $1\%$ 92% $0\%$ $69\%$ $0\%$ $19\%$ 27% 21%
PD ASR(.3) 0% 0% 0% 1% 0% 0% 0% 0%
NH ASR(.3) 0% 14% 0% 28% 0% 2% 0% 2%
[19] PH ASR(.3)$=$ 90%, ($L_{2}=0.2$, HSJA with $3\times 10^{4}$ queries)
[17] PD ASR(.3)$=$ 73%, ($L_{2}=0.1$, up to $8\times 10^{6}$ queries)
4.3 Experiment on Hash-Inversion Attack and Defense
Experiments were conducted over three datasets: MNIST, CelebA, and STL-10. For each hashing algorithm and dataset, we trained the inversion network depicted in Fig. 1 for 50 epochs, with a batch size of 64 and an initial learning rate of 0.005. We utilized the MSE loss function with the AdamW optimizer and the cosine annealing learning rate scheduler. While MNIST and CelebA models were relatively easy to train, we encountered substantial challenges when training STL-10 models. We had to degrade STL-10 images to black-and-white in order to get meaningful reconstructed images.
Some sample images are shown in Fig. 2 whereas major experiment results are left to Appendix A.5 to save space. In general, hash-inversion failed over STL-10 while some level of success could be seen in MNIST and CelebA over PhotoDNA hash only. The reason is that the MNIST and CelebA images have similar or regular formation. The reconstructed images tend to converge to such common formation when the hash is long enough. However, even in this case, the reconstructed CelebA images could not be used to recognize the original face.
The effectiveness of random hash perturbation defense can be clearly seen in Fig. 2(b). Even the slightest perturbation $q=0.1$ made the reconstructed image quite different from the original image.
(a) STL-10 image samples with NeuralHash, without defense
(b) CelebA image samples with PhotoDNA, under various levels $q$ of defense
Figure 2: Samples of true/original images and reconstructed images by the hash inversion algorithm.
5 Conclusions
This paper evaluates the security of three widely used perceptual hashing algorithms (PHAs): PhotoDNA, PDQ, and NeuralHash. Contrary to many existing studies, our findings show that these PHAs demonstrate notable robustness against both adversarial blackbox hash-evasion attacks and hash-inversion attacks. We attribute this robustness, in part, to the random variations inherent in the hash generation process. Building on this insight, we propose enhancing security by introducing additional randomization to the hash bits.
Appendix A Appendix / supplemental material
A.1 Proof of Proposition 1 of Section 3.3
@@ Appendix A Appendix / supplemental material
Without loss of generality, consider the optimization problem (1). A common approach involves utilizing $K$ randomly perturbed images to query the PHAs and then using the queried hashes to estimate the gradient as follows
@@
A.2 Robustness Problem of PHAs to Normal Image Editing
One of the special issue of PHAs is that they seem not robust to some normal imaging editing methods, for which [32] examined PhotoDNA, while [11] investigated PDQ and NeuralHash. Both studies demonstrated that PHAs were robust to certain edits like compression but vulnerable to others like cropping. While such prior works have already extensively investigated image editing on PHAs, a new evaluation is necessary due to ongoing algorithm updates. Notably, NeuralHash has received at least one update since its first release. This study addresses the gap in knowledge by evaluating the robustness of the updated algorithms and, more importantly, by providing the first comparison of the three PHAs.
@@ A.2 Robustness Problem of PHAs to Normal Image Editing
Experiment results are presented in Tables 3. The data for ASR(.3) is highlighted in black because $p=0.3$ is the threshold we suggested to focus on. It is evident that all three PHAs were robust to compression and resizing, but not to filtering and rotation. In conjunction with the findings in [11][16], the PHAs are robust to many standard image editing operations except the following:
@@
NeuralHash was the most robust among the three. Surprisingly, the newest edition of NeuralHash that we experimented with was still not robust to filtering, rotating, and mirroring, similar to the first edition experimented in [11][16]. Nevertheless, we believe this robustness problem can be mitigated by training the model to be invariant to such standard operations with data augmentations.
Table 3: Adversary’s ASR($p$) with Image Editing Attacks
@@
A.3 Extra Experiment Data of Targeted Hash-Evasion Attacks in Section 4.2
Besides the data shown in Section 4.2 and Table 1, more experiment data of targeted attacks are presented in Table 4. We find that the robustness is extremely strong because of ASR$(p)\approx 0$ across all $p$. The only exception was NeuralHash, which exhibited ASR(0.4) of 43% with SimBA. However, this isn’t indicative of successful adversarial attacks; rather, it’s due to numerous hashes generated by NeuralHash having $D_{h}({\bf h}_{0},{\bf h}_{1})$ as low as 0.3. Under NeuralHash, many images are inherently adversarial when evaluated at $p=0.4$.
Therefore, contrary to the prior claims that PHAs were not robust against targeted attacks [19], our findings demonstrate the strong robustness of all PHAs against targeted blackbox attacks. This is significant to the practical application of PHAs because the client adversary needs targeted attacks to ensure that its adversarial hashes do not inadvertently match illicit hashes in the database. The strong robustness means it is hard for the adversary to undertake either untargeted or targeted attacks without being caught.
Table 4: Adversary’s ASR($p$) in Targeted Blackbox Hash-Evasion Attacks. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
ZO- ZOsign
SimBA NES+ sign- -SGD
ASR SimBA +HSJA NES HSJA SGD +HSJA
PH ASR(.1) 0% 0% 0% 0% 0% 0%
PH ASR(.2) 0% 0% 0% 0% 0% 0%
PH ASR(.3) 0% 0% 0% 0% 0% 0%
PH ASR(.4) 1% 2% 0% 0% 0% 0%
PD ASR(.1) 0% 0% 0% 0% 0% 0%
PD ASR(.2) 0% 0% 0% 0% 0% 0%
PD ASR(.3) 0% 0% 0% 0% 0% 0%
PD ASR(.4) 0% 0% 0% 0% 0% 0%
NH ASR(.1) 0% 0% 0% 0% 0% 0%
NH ASR(.2) 0% 0% 0% 0% 0% 0%
NH ASR(.3) 0% 0% 0% 0% 0% 0%
NH ASR(.4) 43% 11% 0% 0% 0% 0%
[19] PH ASR(.3) ($L_{2}=0.19$) 83% ($2\times 10^{6}$ queries)
[19] PD ASR(.3) ($L_{2}=0.42$) 100% ($6\times 10^{5}$ queries)
A.4 Extra Experiment Data of Untargeted Blackbox Hash-Evasion Attacks in Section 4.2
A.4.1 Experiment data and sample images of untargeted attacks without defense
Besides the experiment data shown in Section 4.2 and Table 2, extra experiment data are presented in Table 5. Sample adversarial images generated with our proposed JSHA with various pre-set threshold $p$ are depicted in Fig. 3 and Fig. 4, with their corresponding distortion metrics listed in Table 6. Notably, the images with $p=0.4$ were deemed invalid adversarial images as their distortion was not under the threshold $\theta=0.1$, resulting in extremely noisy images. Similarly, for PDQ, the adversarial images with $p=0.2$ and $0.3$ were also invalid.
Therefore, contrary to previous studies suggesting that PHAs were not robust against untargeted attacks, we found that some PHAs like PDQ and NeuralHash could be considered as robust if realistic constraints regarding image distortion and query budget are taken into account.
Table 5: Adversary’s ASR($p$) in Untargeted Blackbox Hash-Evasion Attacks. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
@@ A.4.1 Experiment data and sample images of untargeted attacks without defense
Table 6: Distortion of the Adversarial Image Samples in Fig. 3(b).
@@
(a) Original image ${\bf I}_{0}$.
PhotoDNA                        PDQ                        NeuralHash
(b) Adversarial images ${\bf I}_{\rm adv}$.
PhotoDNA                        PDQ                        NeuralHash
A.4.2 Experiment data of untargeted attacks under proposed defense
Observing that some untargeted attack algorithms such as SimBA+HSJA achieved relataively high ASR(0.3) over some PHAs such as PhotoDNA, we experimented with the hash perturbation defense method proposed in Section 3.3. The level of our random hash perturbation was denoted as $q$, where $0\leq q\leq 1$, which stand for the ratio of the hash bits that were randomly flipped.
Besides the experiment results presented in Section 4.2 and Table 2, more experiment data are presented in Tables 7, 8, and 9. It can be seen that random perturbation reduced ASR($p$) drastically, which means the attacks became much less successful or even failed. Even a slight perturbation with $q=0.1$ could prevent most of the attacks.
Table 7: Adversary’s ASR($0.1$) in Untargeted Blackbox Hash-Evasion Attacks when Defense was Applied. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
ZOsign
SimBA NES+ -SGD
ASR +HSJA HSJA +HSJA
PH q =0.1 100% $100\%$ $100\%$
PH q =0.2 100% $100\%$ $100\%$
PH q =0.3 100% $100\%$ $100\%$
PD q =0.1 0% 0% 0%
PD q =0.2 0% 0% 0%
PD q =0.3 0% 0% 0%
NH q =0.1 21% 18% 15%
NH q =0.2 20% 18% 15%
NH q =0.3 21% 18% 15%
Table 8: Adversary’s ASR($0.2$) in Untargeted Blackbox Hash-Evasion Attacks when Defense was Applied. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
ZOsign
SimBA NES+ -SGD
ASR +HSJA HSJA +HSJA
PH q =0.1 98% 99% 95%
PH q =0.2 98% 99% 95%
PH q =0.3 98% 99% 95%
PD q =0.1 0% 0% 0%
PD q =0.2 0% 0% 0%
PD q =0.3 0% 0% 0%
NH q =0.1 5% 5% 2%
NH q =0.2 5% 6% 3%
NH q =0.3 5% 6% 2%
Table 9: Adversary’s ASR($0.3$) in Untargeted Blackbox Hash-Evasion Attacks when Defense was Applied. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
ZOsign
SimBA NES+ -SGD
ASR +HSJA HSJA +HSJA
PH q =0.1 27% 21% 5%
PH q =0.2 27% 21% 5%
PH q =0.3 27% 21% 5%
PD q =0.1 0 0% 0%
PD q =0.2 0% 0% 0%
PD q =0.3 0% 0% 0%
NH q =0.1 0% 2% 0%
NH q =0.2 0% 2% 0%
NH q =0.3 0% 0% 0%
A.5 Extra Experiment Results of Hash-Inversion Attacks in Section 4.3